OpenStudy (anonymous):

y = (x^2 + x + 1)/x How to make x the subject of the formula??

mathslover (mathslover):

You mean you want to solve for x ?

mathslover (mathslover):

in terms of y?

OpenStudy (anonymous):

yes.

mathslover (mathslover):

ok so first of all : \[\large{y=\frac{x^2+x+1}{x}}\] \[\large{y*x=x^2+x+1}\]

OpenStudy (anonymous):

"the subject of the formula" what ever happened to "solve for \(x\)"?

mathslover (mathslover):

\[\large{xy=x^2+x+1}\] \[\large{xy-x^2-x=1}\] \[\large{x(y-x-1)=1}\] \[\large{x=\frac{1}{y-x-1}}\]

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

well. that doesn't help. ALL the x terms need to be on one side :P

OpenStudy (anonymous):

I think this is called the inverse of the function.

OpenStudy (anonymous):

you need to solve a quadratic equation

OpenStudy (anonymous):

this is not a one to one function, so doesn't have an inverse, but we can still solve for \(x\)

myininaya (myininaya):

After you get here you wanna put everything on one side and use the quad formula Have fun (hint: group your x terms together) \[\large{xy-x^2-x=1}\]

OpenStudy (anonymous):

\[\large{xy=x^2+x+1}\] \[x^2+x-xy+1=0\] \[x^2+(1-y)x+1=0\] now use the quadatic formula with \(a=1,b=1-y, c=1\) to solve for \(x\)

OpenStudy (anonymous):

See, i don't really need to solve it as much as find an equation for the so called inverse, so that i can find it's domain. the whole point of this is to find the range of the original equation, and to do that, i find the domain of its inverse. if you have any other method, please enlighten me :P

myininaya (myininaya):

Oh Like Sat did.

OpenStudy (anonymous):

yes like i did so you don't have all those pesky negative coefficients!

mathslover (mathslover):

\[\large{x=\frac{1}{y-x-1}}\] \[\large{xy-x^2-x=1}\] \[\large{xy-x^2-x-1=0}\] \[\large{-x^2+x(y-1)-1=0}\] \[\large{x=\frac{1-y\pm \sqrt{(y-1)^2-4}}{-2}}\] \[\large{x=\frac{1-y\pm \sqrt{y^2+1-2y-4}}{-2}}\]

OpenStudy (anonymous):

no this is the method you need so use solve for \(x\) and then note that the discrinant (the expression inside the radical) must be greater than or equal to zero solve that to get your range

OpenStudy (anonymous):

*discriminant

OpenStudy (anonymous):

why greater than or equal to 0, though?? it has 2 distinct real roots??

OpenStudy (anonymous):

you cannot take the square root of negative number

OpenStudy (anonymous):

DUH.

OpenStudy (anonymous):

so in order for \(x\) to be a real number, it must be the case that the expression inside the radical has to be non-negative

OpenStudy (anonymous):

OH.

OpenStudy (anonymous):

assuming that there is no mistake in the work done by mathlover above, the expression inside the radical is \((y-1)^2-4\) so you have to solve \((y-1)^2-4\geq 0\) for \(y\) this will give you the range

OpenStudy (anonymous):

away from this topic, though, for one sec, did you know you had 1287 fans??? THATS INSANE.

OpenStudy (anonymous):

yes, many of them are from that category i assume

OpenStudy (anonymous):

But anywayss, THANK YOU so muchh :D

OpenStudy (anonymous):

i hope it is clear why this works you are looking for the range of \[f(x)=\frac{x^2+x+1}{x}\] so you find the "inverse" and this means what? you pick a \(y\) and say "find the \(x\) that produces it if you pick a \(y\) with \((y-1)^2-4<0\) there will be no such \(x\) whereas if you pick a \(y\) with \[(y-1)^2-4\geq 0\] there will be such an \(x\) that is why this will give the range

OpenStudy (anonymous):

a quick check shows @amistre64 has more fans than i do, leading me to believe that his asylum base is even larger than mine

OpenStudy (amistre64):

i have fans?

OpenStudy (anonymous):

paper ones, imported from taiwan

OpenStudy (amistre64):

was you around when fan count was the equivalent of a "medal" count/smartscore?

OpenStudy (anonymous):

*